# Uncertainty-Aware PPG-2-ECG for Enhanced Cardiovascular Diagnosis using Diffusion Models

### v1

@@ 5.1 Diffusion-Based Conversion: Quality Assessment
Baseline Methods CardioGAN [13] 0.54 99.0 -
RDDM [22] (T=50) 0.22 6.71 -
@@
Table 1: MIMIC-III Results: Quality assessment of our diffusion-based conversion model compared to state-of-the-art baseline methods: CardioGAN [13] and RDDM [22]. The results show the mean metric estimates and their standard errors across the three seeds.
We begin by assessing the quality of our diffusion-based PPG-2-ECG conversion model using the MIMIC-III [38] dataset, showcasing its superiority over state-of-the-art baseline methods. Here, we examine the quality of $\{\hat{X}_{i}\}_{i=1}^{K}$ with $K=100$, sampled from $\hat{\pi}(X|Y)$.
Table 1 presents the conversion performance results, evaluating the quality of the obtained ECG signals. we provide a comparison with two state-of-the-art methods, CardioGAN [13] and RDDM [22], using their reported metrics. More specifically, we assess the conversion quality through RMSE calculations between the generated ECG(s) and the ground truth signals. Additionally, we employ the Fréchet Distance (FD) to evaluate the perceptual quality of the produced signals, considering either a single ECG sample from each PPG (1-FD) or 100 samples from each (100-FD). The results in Table 1 demonstrate the superiority of our conversion model. Observe that by averaging the $K=100$ samples, we obtain an approximation of the MMSE estimator, yielding a comparable performance to the reported result in [22]. When considering the individual ECG solutions for each PPG, the RMSE values increases, as expected [36]. In terms of perceptual quality, the single-solution and multi-solution approaches show pronounced improvement in FD compared to the baselines. Note that computing the FD on multiple solutions helps in improving further the signals’ quality.
@@ 5.2 Classifying Cardiovascular Conditions
To demonstrate the benefits of our diffusion-based synthesized ECGs, we present classification results on two datasets, MIMIC-III [38] and CinC [35], analyzing the classification performance across a series of 11 cardiovascular conditions mentioned earlier. For both datasets, we compare our approach to classification strategies utilized by prior work, varied by the choice of $\pi(Z^{\prime})$ in Definition 3.1, which refers to as the SSC. Specifically, we examine the classification performance of the MMSE estimate and a random single sample from the approximated posterior distribution, $\hat{\pi}(X|Y)$. For more details on the classification strategies considered, we refer the reader to Appendix F.
Note that MIMIC-III lacks cardiovascular labels, and thus cannot be utilized as-is for classification. As a remedy, we apply our experiments using the CinC dataset. However, while CinC contains ECG signals (all leads) and their corresponding labels, it does not include PPG signals, and as such it cannot serve our overall PPG classification goals. We resolve this difficulty by generating synthetic PPG signals from the given ECG ones, using a reversed diffusion-based methodology. The idea is to sample valid and random PPG signals that correspond to their ECG (lead II) counterparts, this way augmenting CinC to form a complete suite. In Appendix G we provide empirical evidence for the validity of our reversal conversion model of ECG-2-PPG in comparison with the original PPG signal.
We now turn to describe the main classification results of cardiovascular conditions, referring to the augmented CinC dataset, and using the synthesized PPG signals. Figure 3 presents ROC curves for the classification strategies considered, and Figure 4 completes this description by presenting risk-coverage curves [25], which evaluate the performance of different confidence functions, denoted as $\kappa$, for PPG signals in regard to uncertainty quantification. For detailed numerical results, including standard errors, refer to Appendix H.
As can be seen from these two figures, our approach (GREEN) demonstrates superior performance over all the alternatives, apart from the direct ECG (BLUE) classification, which serves as an upper-bound for the achievable performance.
Theorem A.1 may be interpreted as claiming that our approach is expected to align with that of classifying the original PPG signals directly (ORANGE), under optimal conditions. However, we observe that our approach achieves better performance. We argue that this gap is due to the erroneous labels used in the direct classification: Note that these labels refer to the original ECG signals, and do not necessarily match the corresponding PPG, which may correspond to other valid ECG signals.
@@
Figure 3: CinC Results: ROC curves for the classification performance of various strategies (see Appendix F), showcasing the superiority of our approach. Higher curves indicate better performance.
Figure 4: CinC Results: Risk-Coverage curves [25] for the uncertainty quantification performance of various strategies, showcasing the superiority of our approach. Lower is better performance.
@@ 6 Concluding Remarks
Our work presents rich empirical results showcasing the superior classification performance of our approach, compared to direct classification methods operating on given PPG signals. Nevertheless, these results are partially reliant on synthetic PPG signals. While we provide empirical validation for these synthetic signals in this paper, further investigation should be dedicated for the study of our approach with original PPG signals, along with further exploration of additional cardiovascular conditions. Future studies in these areas will require additional data collection.
@@ A.1 Optimal Settings
We begin by showing the optimality of the ESC in optimal settings:
Theorem A.1 (Optimality of the Expected Score Classifier).
  1. 1.
  2. 2.
The following is the proof:
Consider the Markovian dependency chain $Y\rightarrow X\rightarrow C$, implying that $\pi(C|X,Y)=\pi(C|X)$, and recalling the assumptions of Theorem A.1:
@@ A.2 Optimality for General Classifiers
Appendix B Assessing Reliable PPG Signals via Calibration
@@
The calibration algorithm is based on a conformal prediction scheme [39, 40, 41], a general method for distribution-free risk control. Specifically, we employ the Learn Then Test (LTT) [42] procedure in this paper. The conformal prediction’s objective is to bound a user-defined risk function, denoted as $R(\lambda)\in(0,1)$, by adjusting a calibration parameter $\lambda\in\mathbb{R}^{+}$ based on empirical data, aiming for the tightest upper bound feasible [43]. Through this calibration process, our aim is to identify $\hat{\lambda}$ that ensures the following guarantee:
@@
where $r_{\delta}\in(0,1)$ serves as a concentration bound, such as Hoeffding’s bound, which is calculated as $r_{\delta}=\sqrt{\frac{1}{2m}\log\frac{1}{\delta}}$. It is shown in [44] that $\mathbb{P}(\hat{R}^{+}(\lambda)<R(\lambda))<\delta$.
@@
Appendix C Conversion-Classification Uncertainty Evidence
Conversion Uncertainty                                                            Classification Uncertainty
Figure 5: CinC Results: Empirical evidence of inherent uncertainty in the PPG-2-ECG conversion-classification framework. (a) The conversion spread and variability as proposed by [24]. High reconstruction error with many PCs indicates high uncertainty in terms of spread and variability. (b) Classification uncertainty shows relatively small interval sizes of ECG probability scores for each PPG (right), which can be effectively modeled using just 100 ECG samples (left).
Here, we present empirical evidence of the inherent uncertainty in the proposed PPG-2-ECG conversion, along with the classification uncertainty implied. Figure 5 summarizes the results that quantify these uncertainty layers, referring to the CinC [35] dataset.
In Figure 5(a), we illustrate the uncertainty as proposed by [24], showing the maximal reconstruction error of the projected ground truth ECG signals among $90\%$ of the temporal signal values using a varying number of Principal Components (PCs) of the ECG solutions’ cloud. Intuitively, a high reconstruction error using many PCs indicates high uncertainty in terms of the spread and variability of possible solutions to the inverse problem tackled. For additional information and details on this metric, we refer the reader to previous work [24].
The above takes us to the results shown in Figure 5(b), depicting the uncertainty in classification, as arises from potential variability in the conversion process. Inspired by [23], in Figure 5(b) (right) we present histograms of interval sizes that contain 90% of the probability scores extracted using the ECG cloud that corresponds to each PPG signal, denoted as $\{f_{\mathcal{X}}(\hat{X}_{i})\}_{i=1}^{K}$. These histograms refer to the series of 11 cardiovascular conditions. Small intervals in this context correspond to a high certainty in classification, and the results reveal that for most of the conditions assessed this is indeed the case. We also see that a small fraction of PPG signals may exhibit a wide range of possible probabilities among ECG candidates, a fact that may motivate our later selective classification approach.
Appendix D Quality Assessment of ECG Visualization Approaches
Signal Domain Embedding Domain
Visualization Strategy RMSE ↓ 1-FD ↓ RMSE ↓ 1-FD ↓
Appendix E Experimental Details
@@
E.1 Datasets and Preprocessing
In this work we rely on two datasets: the MIMIC-III matched waveform database [38], which consists of pairs of measured ECG signals, denoted as $X$, and corresponding PPG signals, denoted as $Y$; and the “Computing in Cardiology” dataset (CinC) [35], which contains pairs of measured ECG signals along with associated cardiovascular condition labels, referred to as $C$.
@@ E.1 Datasets and Preprocessing
    Signal Smoothing: A Butterworth bandpass filter is applied to all signals to reduce noise.
@@
Moving to CinC, which to the best of our knowledge, is the largest and most challenging classification dataset in cardiology, containing 94,630 of 10-second multi-labeled ECG recordings from distinct patients. Each recording is labeled with the presence or absence (binary label) of 11 cardiovascular conditions: (1) Ventricular premature beats, (2) Atrial flutter, (3) Sinus rhythm, (4) Atrial fibrillation, (5) Supraventricular premature beats, (6) Bradycardia, (7) Pacing rhythm, (8) Sinus tachycardia, (9) 1st degree AV block, (10) Sinus bradycardia, and (11) Sinus arrhythmia. We note that the labels considered in CinC exhibit significant imbalance.
@@
    Signal Smoothing: A Butterworth bandpass filter is employed to smooth the ECG signals.
@@
E.2 Training Details for the Conversion Model
Our study leverages a diffusion-based conversion methodology. Specifically, we adopt the DDIM framework [28]. All the training details for the reversed conversion model, designed for generating PPG signals given ECG ones, are similar to the training details for the ECG-2-PPG conversion, as described hereafter.
The denoising model employed in our experiments is based on a U-Net [45] architecture with an attention mechanism, as proposed in [27, 29]. Originally introduced in the field of image processing, this architecture comprises a series of residual layers and downsampling convolutions, followed by a subsequent series of residual layers with upsampling convolutions. Additionally, a global attention layer with a single head is incorporated, along with a projection of the timestep embedding into each residual block. To adapt the architecture for our purposes, we replace the convolutions and attention layers with 1D counterparts, enabling training and evaluation directly in the raw 1D signal domain. In our study we use $d=1024$ PPG and ECG signal length.

### v3

@@ 5.1 Diffusion-Based Conversion: Quality Assessment
Baseline Methods CardioGAN 0.54 99.0 -
RDDM (T=50) 0.22 6.71 -
@@
Table 1: MIMIC-III Results: Quality assessment of our diffusion-based conversion model compared to state-of-the-art baseline methods: CardioGAN (Sarkar & Etemad, 2021) and RDDM (Shome et al., 2023). The results show the mean metric estimates and their standard errors across the three seeds.
We begin by assessing the quality of our diffusion-based PPG-2-ECG conversion model using the MIMIC-III (Johnson et al., 2016) dataset, showcasing its superiority over state-of-the-art baseline methods. Here, we examine the quality of $\{\hat{X}_{i}\}_{i=1}^{K}$ with $K=100$, sampled from $\hat{\pi}(X|Y)$.
Table 1 presents the conversion performance results, evaluating the quality of the obtained ECG signals. we provide a comparison with two state-of-the-art methods, CardioGAN (Sarkar & Etemad, 2021) and RDDM (Shome et al., 2023), using their reported metrics. More specifically, we assess the conversion quality through RMSE calculations between the generated ECG(s) and the ground truth signals. Additionally, we employ the Fréchet Distance (FD) to evaluate the perceptual quality of the produced signals, considering either a single ECG sample from each PPG (1-FD) or 100 samples from each (100-FD). The results in Table 1 demonstrate the superiority of our conversion model. Observe that by averaging the $K=100$ samples, we obtain an approximation of the MMSE estimator, yielding a comparable performance to the reported result in Shome et al. (2023). When considering the individual ECG solutions for each PPG, the RMSE values increases, as expected (Blau & Michaeli, 2018). In terms of perceptual quality, the single-solution and multi-solution approaches show pronounced improvement in FD compared to the baselines. Note that computing the FD on multiple solutions helps in improving further the signals’ quality.
@@ 5.2 Classifying Cardiovascular Conditions
Figure 3: CinC Results: ROC curves for the classification performance of various strategies (see Appendix E). Higher curves indicate better performance.
Figure 4: CinC Results: Risk-Coverage curves for the uncertainty quantification performance of various strategies (see Appendix E). Lower curves indicate better performance.
To demonstrate the benefits of our diffusion-based synthesized ECGs, we present classification results on two datasets, MIMIC-III (Johnson et al., 2016) and CinC (Reyna et al., 2021), analyzing the classification performance across a series of 11 cardiovascular conditions mentioned earlier. For both datasets, we compare our approach to classification strategies utilized by prior work, varied by the choice of $\pi(Z^{\prime})$ in Definition 3.1, which refers to as the SSC. Specifically, we examine the classification performance of the MMSE estimate and a random single sample from the approximated posterior distribution, $\hat{\pi}(X|Y)$. For more details on the classification strategies considered, refer to Appendix E.
Note that MIMIC-III lacks cardiovascular labels, and thus cannot be utilized as-is for classification. As a remedy, we apply our experiments using the CinC dataset. However, while CinC contains ECG signals (all leads) and their corresponding labels, it does not include PPG signals, and as such it cannot serve our overall PPG classification goals. We resolve this difficulty by generating synthetic PPG signals from the given ECG ones, using a reversed diffusion-based methodology. The idea is to sample valid and random PPG signals that correspond to their ECG (lead II) counterparts, this way augmenting CinC to form a complete suite. In Appendix F we provide empirical evidence for the validity of our reversal conversion model of ECG-2-PPG in comparison with the original PPG signal.
We now turn to describe the main classification results of cardiovascular conditions, referring to the augmented CinC dataset, and using the synthesized PPG signals. Figure 3 presents ROC curves for the classification strategies considered, and Figure 4 completes this description by presenting risk-coverage curves (Geifman & El-Yaniv, 2017), which evaluate the performance of different confidence functions, denoted as $\kappa$, for PPG signals in regard to uncertainty quantification. For supplementary results, refer to Appendix G.
As can be seen from these two figures, our approach (GREEN) demonstrates superior performance over all the alternatives, apart from the direct ECG (BLUE) classification, which serves as an upper-bound for the achievable performance. In Appendix H, we provide a justification for this property.
Theorem 3.1 may be interpreted as claiming that our approach is expected to align with that of classifying the original PPG signals directly (ORANGE), under optimal conditions. However, we observe that our approach achieves better performance. We argue that this gap is due to the erroneous labels used in the direct classification: Note that these labels refer to the original ECG signals, and do not necessarily match the corresponding PPG, which may correspond to other valid ECG signals.
@@
5.3 Which ECG Solution to Deliver to Medical Professionals?
Signal Domain Embedding Domain
Visualization Strategy RMSE ↓ 1-FD ↓ RMSE ↓ 1-FD ↓
@@ 6 Concluding Remarks
Our work presents rich empirical results, demonstrating the superior classification performance of our approach compared to direct classification methods operating on given PPG signals. However, these results are partially reliant on synthetic PPG signals, for which we provide empirical validation, though further exploration is required to study additional cardiovascular conditions and the potential risk of hallucinations in the signal domain. Additionally, although the motivation for using PPG-2-ECG is based on the widespread use of wearable devices, no database containing signals from wearable settings was utilized due to the lack of publicly available data; instead, data from hospital facilities were used. Future studies in these areas will require additional data collection.
@@ A.1 Optimal Settings
We begin by presenting a proof for Theorem 3.1 in optimal settings:
Consider the Markovian dependency chain $Y\rightarrow X\rightarrow C$, implying that $\pi(C|X,Y)=\pi(C|X)$, and recalling the assumptions of Theorem 3.1:
@@ A.2 Optimality for General Classifiers
Appendix B Uncertainty Quantification
The proposed PPG-2-ECG conversion has two layers of uncertainty awareness, as described herein. The first takes into account the inherent uncertainty that emerges from the ill-posed conversion itself, quantifying the spread and variability of the possible ECG candidates that derive from the PPG inputs. In Appendix B.1, we present empirical evidence illustrating the uncertainty from this perspective. The second uncertainty layer leverages the above property for providing classification predictions with a better confidence in their correctness. This serves a selective classification (Geifman & El-Yaniv, 2017) scheme, in which a carefully chosen subset of the test PPG data is selected by their calculated confidence, exhibiting improved accuracy.
Selective classification (Geifman & El-Yaniv, 2017) enables handling cases in which classifying every PPG signal within a test set is challenging due to uncertainty or data corruption. This technique allows for abstaining from making predictions for signals with low confidence, thus reducing classification errors on the remaining data. A confidence function, defined as $\kappa:\mathcal{Y}\rightarrow\mathbb{R}^{+}$, serves as the engine for this task by predicting the correctness of predictions. One common choice of this confidence function $\kappa$ is to use the maximal softmax score of a classifier. In a similar spirit, in this work we set $\kappa(Y)$ to take the value $\kappa(Y)=\max\{1-f_{\mathcal{Y}}(Y),f_{\mathcal{Y}}(Y)\}$, where $f_{\mathcal{Y}}(Y)$ approximates the ESC (Definition 3.2), as summarized in Algorithm 1.
where $\delta\in(0,1)\ll 1$ is a user-defined calibration error. Intuitively, by setting $\alpha=\delta=0.1$, the above implies that the classification error among selected PPG signals, whose confidence scores are greater than $\lambda$, will be lower than $\alpha$ ($=10\%$) with probability of $1-\delta$ ($=90\%$). In Appendix B.2, we describe how to calibrate $\lambda$ to provide the guarantee stated in Equation (5).
B.1 Conversion-Classification Uncertainty Evidence
Conversion Uncertainty                                                 Classification Uncertainty
Figure 5: CinC Results: Empirical evidence of inherent uncertainty in the PPG-2-ECG conversion-classification framework. (a) The conversion spread and variability as proposed by Belhasin et al. (2023). High reconstruction error with many PCs indicates high uncertainty in terms of spread and variability. (b) Classification uncertainty shows relatively small interval sizes of ECG probability scores for each PPG (right), which can be effectively modeled using just 100 ECG samples (left).
Here, we present empirical evidence of the inherent uncertainty in the proposed PPG-2-ECG conversion, along with the classification uncertainty implied. Figure 5 summarizes the results that quantify these uncertainty layers, referring to the CinC (Reyna et al., 2021) dataset.
In Figure 5(a), we illustrate the uncertainty as proposed by Belhasin et al. (2023), showing the maximal reconstruction error of the projected ground truth ECG signals among $90\%$ of the temporal signal values using a varying number of Principal Components (PCs) of the ECG solutions’ cloud. Intuitively, a high reconstruction error using many PCs indicates high uncertainty in terms of the spread and variability of possible solutions to the inverse problem tackled. For additional information and details on this metric, we refer the reader to previous work by Belhasin et al. (2023).
The above takes us to the results shown in Figure 5(b), depicting the uncertainty in classification, as arises from potential variability in the conversion process. Inspired by Angelopoulos et al. (2022), in Figure 5(b) (right) we present histograms of interval sizes that contain 90% of the probability scores extracted using the ECG cloud that corresponds to each PPG signal, denoted as $\{f_{\mathcal{X}}(\hat{X}_{i})\}_{i=1}^{K}$. These histograms refer to the series of 11 cardiovascular conditions. Small intervals in this context correspond to a high certainty in classification, and the results reveal that for most of the conditions assessed this is indeed the case. We also see that a small fraction of PPG signals may exhibit a wide range of possible probabilities among ECG candidates, a fact that may motivate our later selective classification approach.
B.2 Assessing Reliable PPG Signals via Calibration
@@
The calibration algorithm is based on a conformal prediction scheme (Vovk et al., 2005; Papadopoulos et al., 2002; Lei & Wasserman, 2014), a general method for distribution-free risk control. Specifically, we employ the Learn Then Test (LTT) (Angelopoulos et al., 2021) procedure in this paper. The conformal prediction’s objective is to bound a user-defined risk function, denoted as $R(\lambda)\in(0,1)$, by adjusting a calibration parameter $\lambda\in\mathbb{R}^{+}$ based on empirical data, aiming for the tightest upper bound feasible (Angelopoulos & Bates, 2021). Through this calibration process, our aim is to identify $\hat{\lambda}$ that ensures the following guarantee:
@@
where $r_{\delta}\in(0,1)$ serves as a concentration bound, such as Hoeffding’s bound, which is calculated as $r_{\delta}=\sqrt{\frac{1}{2m}\log\frac{1}{\delta}}$. It is shown in Hoeffding (1994) that $\mathbb{P}(\hat{R}^{+}(\lambda)<R(\lambda))<\delta$.
@@
Appendix C Analysis of Overfitting in the Conversion Model
In this section, we provide additional experimental evidence supporting the validity of our conversion model using the MIMIC-III (Johnson et al., 2016) dataset. We analyze the effect of overfitting and demonstrate that our model does not exhibit overfitting.
C.1 Conversion Training Metrics: Training vs. Validation
Figure 6: MIMIC-III Results: Metrics report of MSE loss and PSNR comparing the training set and validation set during training, indicating no overfitting.
Our PPG-2-ECG conversion model is a conditional diffusion model trained on the MIMIC-III (Johnson et al., 2016) dataset. Following the gold standards for training conditional diffusion models, a denoising model is required to clean noise from noisy ECG signals, conditioned on the PPG input.
In Figure 6, we present the average MSE loss and PSNR of the denoised ECG signals compared to the original ones for each epoch during the training process. It is evident that the performance on the training and validation sets is similar, indicating that there is no overfitting. We attribute this to the diversity and large scale of the MIMIC-III dataset.
C.2 Original ECG Frequency Approximation
(a) 25% Quantile Original ECG
(b) Median Original ECG
(c) 75% Quantile Original ECG
Figure 7: MIMIC-III Results: Pairwise synthetic ECG distances (GREEN) for three PPG signals and the nearest synthetic ECG distance to the original ECG (BLUE). Results show the original ECG’s approximation within the synthetic solution cloud across three distance quantiles of each original ECG.
Here, we present an additional signal-quality assessment, demonstrating how often the original ECG can be approximated as one of the synthetic ECG signals within each solution cloud for each PPG.
To evaluate this, we first calculated histograms of pairwise euclidean distances between synthetic ECG signals within the solution cloud for each PPG signal in the test set. Next, we determined the distance from the original ECG to its nearest synthetic ECG signal. Then, we assessed how often the distance of the original ECG fell within the 99% of the histogram. We found that 93.23% of the original ECG signals were contained within the solution clouds across the test set, indicating that our solution cloud provides a suited approximation of the original ECG signals.
In Figure 7, we visually demonstrate the inclusion of the original ECG distance within three histograms of three different PPG signals.
Appendix D Experimental Details
@@
D.1 Datasets and Preprocessing
In this work we rely on two datasets: the MIMIC-III matched waveform database (Johnson et al., 2016), which consists of pairs of measured ECG signals, denoted as $X$, and corresponding PPG signals, denoted as $Y$; and the “Computing in Cardiology” dataset (CinC) (Reyna et al., 2021), which contains pairs of measured ECG signals along with associated cardiovascular condition labels, referred to as $C$.
D.1.1 MIMIC-III Dataset Summary
@@
    Signal Smoothing: A zero-phase 3rd order Butterworth bandpass filter with a frequency band ranging from 1 Hz to 47 Hz is applied to all signals to reduce noise.
@@
D.1.2 CinC Dataset Summary
Figure 8: CinC Results: Class distribution report, indicating a significant class imbalance.
Moving to CinC, which to the best of our knowledge, is the largest and most challenging classification dataset in cardiology, containing 88,253 of 10-second multi-labeled ECG recordings from distinct patients. CinC integrates data from several public databases, including the CPSC Database and CPSC-Extra Database, INCART Database, PTB and PTB-XL Database, the Georgia 12-lead ECG Challenge (G12EC) Database, Augmented Undisclosed Database, Chapman-Shaoxing and Ningbo Database, and the University of Michigan (UMich) Database. For further details on these databases, we refer the reader to Reyna et al. (2021).
Each recording is labeled with the presence or absence (binary label) of various cardiovascular conditions. From these, 11 conditions detectable using lead II ECG alone were selected: (1) Ventricular premature beats, (2) Atrial flutter, (3) Sinus rhythm, (4) Atrial fibrillation, (5) Supraventricular premature beats, (6) Bradycardia, (7) Pacing rhythm, (8) Sinus tachycardia, (9) 1st degree AV block, (10) Sinus bradycardia, and (11) Sinus arrhythmia. We note that the labels considered in CinC exhibit significant imbalance, as presented in Figure 8.
@@ D.1.2 CinC Dataset Summary
    Signal Smoothing: A zero-phase 3rd order Butterworth bandpass filter with a frequency band ranging from 1 Hz to 47 Hz is employed to smooth the ECG signals.
@@
D.2 Training Details for the Conversion Model
Our study leverages a diffusion-based conversion methodology. Specifically, we adopt the DDIM framework (Song et al., 2020). All the training details for the reversed conversion model, designed for generating PPG signals given ECG ones, are similar to the training details for the ECG-2-PPG conversion, as described hereafter.
The denoising model employed in our experiments is based on a U-Net (Ronneberger et al., 2015) architecture with an attention mechanism, as proposed in Ho et al. (2020); Dhariwal & Nichol (2021). Originally introduced in the field of image processing, this architecture comprises a series of residual layers and downsampling convolutions, followed by a subsequent series of residual layers with upsampling convolutions. Additionally, a global attention layer with a single head is incorporated, along with a projection of the timestep embedding into each residual block. To adapt the architecture for our purposes, we replace the convolutions and attention layers with 1D counterparts, enabling training and evaluation directly in the raw 1D signal domain. In our study we use $d=1024$ PPG and ECG signal length.